Comey Products has decided to acquire some new equipment having a $260,000 purchase price. The equipment will last 4 years and is in the MACRS 3-year class. (The depreciation rates for Year 1 through Year 4 are equal to 0.3333, 0.4445, 0.1481, and 0.0741.) The firm can borrow at a 10% rate and pays a 25% federal-plus-state tax rate. Comey is considering leasing the property but wishes to know the cost of borrowing that it should use when comparing purchasing to leasing and has hired you to answer this question. What is the correct answer to Comey’s question? (Hint: Use the shortcut method to find the after-tax cost of the loan payments.) Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

After tax cost of debt = I% * (1-T%). Here I% = borrowing rate= 10% , T%= Tax rate of the company= 25%

Thus after tax cost of debt = 0.1 * (1-0.25) = 0.1 * 0.75 = 0.075 = 7.50%

7.50% is the after tax cost of borrowing which Comey Products should use when comparing purchasing to leasing.